# A new infinitesimal form of the Prékopa–Leindler inequality with multiplicative structure and applications

### v1

@@ 1 Introduction
A deeper consequence of the log-concavity property (2) is the validity of certain Poincaré-type inequalities on the boundaries of convex sets. As before, suppose that $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$ where $u:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ is a lower semi-continuous convex function. Assume in addition that $u\in C^{1}(\mathbb{R}^{n})$ and consider a compact convex set $K\subset\mathbb{R}^{n}$ whose boundary $\partial K$ is a strictly convex $C^{2}$ hypersurface. Then, for every locally Lipschitz function $\rho:\partial K\to\mathbb{R}$, the following inequality holds:
@@
Here, $\rm II$} denotes the second fundamental form of $\partial K$, $\nabla_{\partial K}$ the tangential gradient, and $\langle\cdot,\cdot\rangle$ the restriction of the Euclidean scalar product to the tangent bundle of $\partial K$. With a slight abuse of notation, we use the same symbol for the scalar product on the whole space. The weighted mean curvature is given by ${\rm H}_{\mu}=\mathrm{tr}({\rm II})-\langle\nabla u,\nu_{K}\rangle$, where $\nu_{K}$ denotes the outer unit normal to $\partial K$. Integration on $\partial K$ with respect to $\mu$ is understood as integration against the (n1)(n-1)-dimensional Hausdorff measure weighted by $e^{-u}$. An equivalent form of (3) was proved by Colesanti [11] in the case of the Lebesgue measure, and extended to a much more general setting by Kolesnikov and Milman [25].
@@
Statements of this kind are known as concavity principles for marginals. For more details on the topic, see, for example, the recent work of Cordero-Erasquin and Eskenazis [14], which was of great inspiration in the preparation of this manuscript. If one proves that $\mathcal{S}$ is twice differentiable as $t\to 0^{+}$, the concavity of $\mathcal{S}$ can be equivalently stated as $\mathcal{S}^{\prime\prime}(0)\leq 0$. Writing explicitly $\mathcal{S}^{\prime\prime}(0)$ gives, for appropriate choices of $\Phi$, the inequalities (3) and (4).
@@
The Poincaré inequality (3) and the Brascamp–Lieb inequality (4), together with suitable variants, have been the subject of intensive research in the last couple of decades. As mentioned above, Colesanti [11] obtained a non-weighted version of (3). In the same period, Bobkov and Ledoux [3, 4] generalized (4) to $\beta$-concave functions computing an infinitesimal form of the Borell–Brascamp–Lieb inequality (compare [5, 9]), a generalization of the Prékopa–Leindler inequality (1). We point out that our methods can be easily extended to obtain a generalization of Theorem 1.1 stemming from the Borell–Brascamp–Lieb inequality. Nonetheless, we decided to focus on the log-concave case for a cleaner and more accessible exposition. Around the same time, Cordero-Erasquin, Fradelizi, and Maurey [13] used (4) to prove the (B)-conjecture for the Gaussian measure, pioneering the so-called $L_{2}$ method in the context of geometric and analytic inequalities. Almost a decade later, Kolesnikov and Milman [24, 25] extended this approach to the setting of weighted Riemannian manifolds under suitable curvature assumptions, reigniting interest in the topic. With similar techniques, Nguyen [33] extended the results of [3]. Afterwards, the application of the $L_{2}$ method has found several applications in the extension of the (B)-conjecture to even log-concave probability measures [14, 15, 16] (which we will not treat in this work) and to the dimensional Brunn–Minkowski inequality [14, 16, 17, 22, 23, 29] (which we present later in (7)). Both conjectures, already interesting on their own, have acquired further relevance in relation to the logarithmic Brunn–Minkowski conjecture, stated in [7] by Böröczky, Lutwak, Yang, and Zhang, and are intimately connected to the logarithmic Minkowski problem, introduced by the same authors in [8]. It was proved by Saroglou [37] that the log-Brunn–Minkowski conjecture implies the dimensional Brunn–Minkowski inequality, and by Livshyts, Marsiglietti, Nayar, and Zvavitch [28] that a solution of the log-Brunn–Minkowski conjecture would imply the (B)-conjecture. Further details on the development of these open problems can be found in [14].
@@
Let us now shortly present the dimensional Brunn–Minkowski conjecture. We say that $A\subset\mathbb{R}^{n}$ is origin-symmetric if $A=-A=\{-x\colon x\in A\}$, and that a measure $\mu$ is even if $\mu(B)=\mu(-B)$ for every $\mu$-measurable set $B\subset\mathbb{R}^{n}$. The question is the following: Given an even log-concave measure $\mu$ on $\mathbb{R}^{n}$, is it true that, for every pair of origin-symmetric compact convex sets $K,L\subset\mathbb{R}^{n}$, the inequality
@@
This conjecture was put forth by Gardner and Zvavitch for the Gaussian measure [20] and was resolved by Eskenazis and Moschidis [17]. Cordero-Erasquin and Rotem [16] gave an affirmative answer for sufficiently regular and rotationally invariant log-concave measures, while Cordero-Erasquin and Eskenazis [14] established the conjecture for hereditarily convex log-concave measures. For general even log-concave measures, Livshyts [29] proved (7) with exponent $n^{-(4+o(1))}$, which currently constitutes the best known universal bound.
@@
It is worth mentioning that the operator $\nabla_{\partial K}\cdot({\rm II}^{-1}\nabla_{\partial K}(\cdot))-\langle\nabla_{\partial K}u,{\rm II}^{-1}\nabla_{\partial K}(\cdot)\rangle$ is nothing but a weighted Laplacian on $\partial K$ with an appropriate choice of metric and measure. Compare, e.g., [25, Section 6.5].
@@
holds, where $\overline{\rho}$ is the function provided by Theorem 1.3. Indeed, $\overline{\rho}$ may be interpreted as an extremal perturbation of $K$, encoding the concavity properties of $K$ while simultaneously incorporating the geometry induced by the measure $\mu$. It is interesting to note that in the presence of symmetry, the minimizer is even. Moreover, we note that estimates similar to (10) were already required in [14] (see Lemma 11 therein) to establish both the (B)-conjecture and the dimensional Brunn–Minkowski conjecture, together with a new spectral inequality. By contrast, the Euler–Lagrange approach adopted here appears to bypass the need for such spectral inequalities, at the cost of a more delicate analysis of the PDE (8).
@@
Consider an origin-symmetric compact convex set $K\subset\mathbb{R}^{n}$ such that $\partial K$ is a strictly convex manifold of class $C^{2}$, and an even strictly convex function $u\in C^{2}(\mathbb{R}^{n})$. Consider the measure $\mu$ with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$ and suppose, moreover, that $k_{1}\mathrm{Id}\leq\nabla^{2}u$ and $\Delta u\leq nk_{2}$ for some $k_{2}\geq k_{1}>0$. Then
@@
where $c=\frac{1}{2(r+1)}$, with $r=\frac{k_{2}}{k_{1}}\geq 1$.
@@
Notice that the estimate obtained in [23] is slightly sharper (even though morally equivalent up to an absolute constant). In the notation above, it corresponds to the constant $c=2/(\sqrt{r}+1)^{2}$. The bound derived here, however, is obtained through a substantially different procedure, building directly on the framework developed in the preceding sections. In particular, we believe this different method to be of independent interest.
@@ 2 A concavity principle along functional Wulff shapes
As we will soon have to compute the first and second derivatives of (13), we need to assume further regularity on the sets and functions involved. We say that $K\in\mathcal{K}^{n}$ is a smooth convex body if it is strictly convex (that is, $\partial K$ does not contain segments) and $\partial K$ is a manifold of class $C^{2}$. In particular, $h_{K}$ is of class $C^{2}$ in $\mathbb{R}^{n}\setminus\{o\}$ (compare [38, Section 2.5]) and the main curvatures of $\partial K$ are strictly positive and bounded as functions on $\partial K$. Additionally, in this section, we will always tacitly assume that $o\in\mathrm{int}K$ (the interior of $K$), as the problems at hand will always be invariant under translation. We say that a function $u\in\mathrm{Conv}(\mathbb{R}^{n})$ is smooth and strictly convex if $u\in C^{2}(\mathbb{R}^{n})$ and $\nabla^{2}u$ is positive definite. We now provide a statement concerning the computation of shape derivatives. In the context we are treating, similar facts have been proved in [25] and [14]. For a function $\Phi_{t}$ differentiable in $t$ (and which might depend on other variables), we adopt the notation $\Phi^{\prime}_{t}=\frac{\partial}{\partial t}\Phi_{t}$ for the partial derivative on the entry of $t$ only.
@@
this vector field is clearly well-defined on $\mathbb{R}^{n}$, Lipschitz, differentiable for every $t\in(-\varepsilon,\varepsilon)$, and, clearly, $X_{0}=\mathrm{Id}$. Finally, to check that $X_{t}$ is invertible for every $t\in(-\varepsilon,\varepsilon)$, observe that, for every $\tau\geq 0$ and every fixed $t\in(-\varepsilon,\varepsilon),$ $X_{t}$ is, by construction, a bijection (in fact, a diffeomorphism) between $\partial(\tau K)$ and $\partial(\tau K_{t})$. Since both $K$ and $K_{t}$ have the origin inside their interior, $\mathbb{R}^{n}$ can be represented as the disjoint unions
@@
$I^{\prime\prime}(0)-\frac{1}{I(0)}I^{\prime}(0)^{2}\leq 0,$ (28)
@@ 3 Stability for the Poincaré inequality
Let $K\subset\mathbb{R}^{n}$ be a smooth convex body and $u:\mathbb{R}^{n}\to\mathbb{R}$ a smooth and strictly convex function, and consider on $K$ the log-concave probability measure $\mu$ with density $\mathrm{d}\mu(x)=\frac{e^{-u(x)}}{\int_{K}e^{-u(y)}\,\mathrm{d}y}\,\mathrm{d}x$. For $\psi\in C^{1}(K)$, consider the weighted Sobolev norm
@@
where $\|x\|^{2}=\langle x,x\rangle$ for every $x\in\mathbb{R}^{n}$. The weighted Sobolev space $H^{1}(K,\mu)$ is defined as the completion of $C^{\infty}(K)$ under $\|\cdot\|_{H^{1}(K,\mu)}$. On $\partial K$ regarded as a weighted (n1)(n{-}1)–dimensional manifold with the measure $\mu_{\partial K}=e^{-u}\,\mathcal{H}^{n-1}$ and tangential gradient $\nabla_{\partial K}$ we set, for $\rho\in C^{1}(\partial K)$,
$\|\rho\|_{H^{1}(\partial K,\mu)}^{2}=\int_{\partial K}\left(\rho^{2}+\|\nabla_{\partial K}\rho\|^{2}\right)\,\,\mathrm{d}\mu_{\partial K}.$
As before, when integrating on the boundary, $\mathrm{d}\mu$ is understood as $\mathrm{d}\mu_{\partial K}$. Again for $\rho\in C^{1}(\partial K)$, we denote its fractional Sobolev norm by
@@
In addition, since $K$ is uniformly convex and compact, the minimal and maximal curvatures of $\partial K$
@@
We can now prove the main result of this section.
@@ 4 Concavity powers of strictly log-concave measures
We now explain how the quantity in (48) is connected to spectral inequalities. Fix a smooth and strictly convex $K\in\mathcal{K}^{n}$, and let $L\in\mathcal{K}^{n}$ be smooth and strictly convex as well. Define the function $\rho:\partial K\to\mathbb{R}$ by
$\rho(\nu_{K}(x))=h_{L}(\nu_{K}(x))-h_{K}(\nu_{K}(x)),$ (49)
@@
$p(\mu,K)=\sup_{p\in(0,\infty)}\left\{p:\left.\frac{d^{2}}{dt^{2}}\mu(K_{t}(\rho))^{p}\right|_{t=0}\leq 0,\ \forall\rho\in C^{1}(\partial K)\right\},$
where we have that linear combinations of differences of support functions in $C^{2}(\mathbb{S}^{n-1})$ are dense in $C^{1}(\mathbb{S}^{n-1})$. If we now write
@@
In the case where the convex sets $K$ and $L$ are origin symmetric, the admissible perturbations $\rho$ in (49) must be even. Consequently, the infimum in (50) is taken over all even functions in $C^{1}(\partial K)$. Using the $L_{2}$ method, the functional $J(\rho)$ can be estimated from below using only integrals over $K$; see, for example, Proposition 3.16 in [30]. This has been the prevailing approach in all prior results in the literature.
@@
The main result of this section reads as follows. We recall that for a vector field $X\in C^{1}(\partial K,\mathbb{R}^{n})$, $\nabla_{\partial K}\cdot X$ is the tangential divergence of $X$.
@@
for every $X\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{n})$.With this notation, we define the operator $\mathcal{L}$ in Theorem 4.3 by
@@
Thus, the function $\overline{\rho}$ is a weak solution of $\mathcal{L}(\overline{\rho})=1$. Notice that, if $K$ is smooth, then the operator $\mathcal{L}$ is uniformly elliptic since ${\rm II}^{-1}$ is positive definite.

### v2

@@ 1 Introduction
A deeper consequence of the log-concavity property (2) is the validity of certain Poincaré-type inequalities on the boundaries of convex sets. As before, suppose that $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$ where $u:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ is a lower semi-continuous convex function. Assume in addition that $u\in C^{1}(\mathbb{R}^{n})$ and consider a compact convex set $K\subset\mathbb{R}^{n}$ whose boundary $\partial K$ is a strictly convex $C^{2}$ hypersurface. We shall say, with a slight abuse of terminology, that $\partial K$ is strictly convex whenever its second fundamental form $\rm II$} is positive definite. Then, for every locally Lipschitz function $\rho:\partial K\to\mathbb{R}$, the following inequality holds:
@@
Here, $\nabla_{\partial K}$ denotes the tangential gradient, and $\langle\cdot,\cdot\rangle$ the restriction of the Euclidean scalar product to the tangent bundle of $\partial K$. With a slight abuse of notation, we use the same symbol for the scalar product on the whole space. The weighted mean curvature is given by ${\rm H}_{\mu}=\mathrm{tr}({\rm II})-\langle\nabla u,\nu_{K}\rangle$, where $\nu_{K}$ denotes the outer unit normal to $\partial K$. Integration on $\partial K$ with respect to $\mu$ is understood as integration against the (n1)(n-1)-dimensional Hausdorff measure weighted by $e^{-u}$. An equivalent form of (3) was proved by Colesanti [11] in the case of the Lebesgue measure, and extended to a much more general setting by Kolesnikov and Milman [25].
@@
Statements of this kind are known as concavity principles for marginals. For more details on the topic, see, for example, the recent work of Cordero-Erausquin and Eskenazis [14], which was of great inspiration in the preparation of this manuscript. If one proves that $\mathcal{S}$ is twice differentiable as $t\to 0^{+}$, the concavity of $\mathcal{S}$ can be equivalently stated as $\mathcal{S}^{\prime\prime}(0)\leq 0$. Writing explicitly $\mathcal{S}^{\prime\prime}(0)$ gives, for appropriate choices of $\Phi$, the inequalities (3) and (4).
@@
The Poincaré inequality (3) and the Brascamp–Lieb inequality (4), together with suitable variants, have been the subject of intensive research in the last couple of decades. As mentioned above, Colesanti [11] obtained a non-weighted version of (3). In the same period, Bobkov and Ledoux [3, 4] generalized (4) to $\beta$-concave functions computing an infinitesimal form of the Borell–Brascamp–Lieb inequality (compare [5, 9]), a generalization of the Prékopa–Leindler inequality (1). We point out that our methods can be easily extended to obtain a generalization of Theorem 1.1 stemming from the Borell–Brascamp–Lieb inequality. Nonetheless, we decided to focus on the log-concave case for a cleaner and more accessible exposition. Around the same time, Cordero-Erausquin, Fradelizi, and Maurey [13] used (4) to prove the (B)-conjecture for the Gaussian measure, pioneering the so-called $L_{2}$ method in the context of geometric and analytic inequalities. Almost a decade later, Kolesnikov and Milman [24, 25] extended this approach to the setting of weighted Riemannian manifolds under suitable curvature assumptions, reigniting interest in the topic. With similar techniques, Nguyen [33] extended the results of [3]. Afterwards, the application of the $L_{2}$ method has found several applications in the extension of the (B)-conjecture to even log-concave probability measures [14, 15, 16] (which we will not treat in this work) and to the dimensional Brunn–Minkowski inequality [14, 16, 17, 22, 23, 29] (which we present later in (7)). Both conjectures, already interesting on their own, have acquired further relevance in relation to the logarithmic Brunn–Minkowski conjecture, stated in [7] by Böröczky, Lutwak, Yang, and Zhang, and are intimately connected to the logarithmic Minkowski problem, introduced by the same authors in [8]. Thanks to the works of Saroglou [37] and Livshyts, Marsiglietti, Nayar, and Zvavitch [28], it is known that the logarithmic Brunn–Minkowski conjecture implies the dimensional Brunn–Minkowski conjecture and the (B)-conjecture. Further details on the development of these open problems can be found in [14].
@@
Let us now shortly present the dimensional Brunn–Minkowski conjecture. We say that $A\subset\mathbb{R}^{n}$ is origin-symmetric if $A=-A=\{-x\colon x\in A\}$, and that a measure $\mu$ is even if $\mu(B)=\mu(-B)$ for every $\mu$-measurable set $B\subset\mathbb{R}^{n}$. The question is the following: Given an even log-concave measure $\mu$ on $\mathbb{R}^{n}$, is it true that, for every pair of origin-symmetric non-empty compact convex sets $K,L\subset\mathbb{R}^{n}$, the inequality
@@
This conjecture was put forth by Gardner and Zvavitch for the Gaussian measure [20] and was resolved by Eskenazis and Moschidis [17]. Cordero-Erausquin and Rotem [16] gave an affirmative answer for sufficiently regular and rotationally invariant log-concave measures, while Kolesnikov and Livshyts [23] established the conjecture for hereditarily convex log-concave measures. For general even log-concave measures, Livshyts [29] proved (7) with exponent $n^{-(4+o(1))}$, which currently constitutes the best known universal bound.
@@
It is worth mentioning that the operator $\nabla_{\partial K}\cdot({\rm II}^{-1}\nabla_{\partial K}(\cdot))+\langle\nabla_{\partial K}u,{\rm II}^{-1}\nabla_{\partial K}(\cdot)\rangle$ is nothing but a weighted Laplacian on $\partial K$ with an appropriate choice of metric and measure. Compare, e.g., [25, Section 6.5].
@@
holds, where $\overline{\rho}$ is the function provided by Theorem 1.3. Indeed, $\overline{\rho}$ may be interpreted as an extremal perturbation of $K$, encoding the concavity properties of $K$ while simultaneously incorporating the geometry induced by the measure $\mu$. It is interesting to note that in the presence of symmetry, the minimizer is even. Moreover, we note that estimates similar to (10) were already required in [14] (see Lemma 11 therein) to establish both the (B)-conjecture and the dimensional Brunn–Minkowski conjecture, together with certain spectral inequalities. By contrast, the Euler–Lagrange approach adopted here appears to bypass the need for such spectral inequalities, at the cost of a more delicate analysis of the PDE (8).
@@
Consider an origin-symmetric compact convex set $K\subset\mathbb{R}^{n}$ such that $\partial K$ is a strictly convex manifold of class $C^{2}$, and an even strictly convex function $u\in C^{2}(\mathbb{R}^{n})$. Consider the measure $\mu$ with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$ and suppose, moreover, that $k_{1}\,\mathrm{Id}\leq\nabla^{2}u$ and $\Delta u\leq k_{2}\,n$ for some $k_{2}\geq k_{1}>0$. Then
@@
where $c=\frac{1}{r+1}$, with $r=\frac{k_{2}}{k_{1}}\geq 1$.
@@
Notice that the estimate obtained in [23] is slightly sharper (even though morally equivalent up to an absolute constant). In the notation above, it corresponds to the constant $c=2/(\sqrt{r}+1)^{2}$. The bound derived here, however, is obtained through a substantially different procedure, building directly on the framework developed in the preceding sections. In particular, we believe this different method to be of independent interest. Finally, we note that, when $\mu$ is the Gaussian measure, Theorem 1.4 implies the main result of [23]; in particular, a close inspection of the proof implies that the symmetry assumption on the convex sets involved can be weakened to the assumption that the sets merely contain the origin.
@@ 2 A concavity principle along functional Wulff shapes
As we will soon have to compute the first and second derivatives of (13), we need to assume further regularity on the sets and functions involved. We say that $K\in\mathcal{K}^{n}$ is a smooth convex body if $\partial K$ is a manifold of class $C^{2}$ and strictly convex in the sense described in the introduction, i.e., ${\rm II}>0$. In particular, $h_{K}$ is of class $C^{2}$ in $\mathbb{R}^{n}\setminus\{o\}$ (compare [38, Section 2.5]) and the main curvatures of $\partial K$ are strictly positive and bounded as functions on $\partial K$. Additionally, in this section, we will always tacitly assume that $o\in\mathrm{int}K$ (the interior of $K$), as the problems at hand will always be invariant under translation. We say that a function $u\in\mathrm{Conv}(\mathbb{R}^{n})$ is smooth and strictly convex if $u\in C^{2}(\mathbb{R}^{n})$ and $\nabla^{2}u$ is positive definite. We now provide a statement concerning the computation of shape derivatives. In the context we are treating, similar facts have been proved in [25] and [14]. For a function $\Phi_{t}$ differentiable in $t$ (and which might depend on other variables), we adopt the notation $\Phi^{\prime}_{t}=\frac{\partial}{\partial t}\Phi_{t}$ for the partial derivative on the entry of $t$ only.
@@
this vector field is clearly well-defined on $\mathbb{R}^{n}$, Lipschitz, differentiable for every $t\in(-\varepsilon,\varepsilon)$, and $X_{0}=\mathrm{Id}$. Finally, to check that $X_{t}$ is invertible for every $t\in(-\varepsilon,\varepsilon)$, observe that, for every $\tau\geq 0$ and every fixed $t\in(-\varepsilon,\varepsilon),$ $X_{t}$ is, by construction, a bijection (in fact, a diffeomorphism) between $\partial(\tau K)$ and $\partial(\tau K_{t})$. Since both $K$ and $K_{t}$ have the origin inside their interior, $\mathbb{R}^{n}$ can be represented as the disjoint unions
@@
$I^{\prime\prime}(0)-\frac{I^{\prime}(0)^{2}}{I(0)}\leq 0,$ (28)
@@ 3 Stability for the Poincaré inequality
Let $K\subset\mathbb{R}^{n}$ be a smooth convex body and $u:\mathbb{R}^{n}\to\mathbb{R}$ a smooth and strictly convex function, and consider on $K$ the log-concave measure $\mu$ with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. For $\psi\in C^{1}(K)$, consider the weighted Sobolev norm
@@
where $\|x\|^{2}=\langle x,x\rangle$ for every $x\in\mathbb{R}^{n}$. The weighted Sobolev space $H^{1}(K,\mu)$ is defined as the completion of $C^{\infty}(K)$ under $\|\cdot\|_{H^{1}(K,\mu)}$. On $\partial K$ we set, for $\rho\in C^{1}(\partial K)$,
$\|\rho\|_{H^{1}(\partial K,\mu)}^{2}=\int_{\partial K}\left(\rho^{2}+\|\nabla_{\partial K}\rho\|^{2}\right)\,\,\mathrm{d}\mu.$
Again for $\rho\in C^{1}(\partial K)$, we denote its fractional Sobolev norm by
@@
In addition, since $K$ is strictly convex and compact, the minimal and maximal curvatures of $\partial K$
@@
We can now prove the main result of this section (Theorem 1.2 in the introduction).
@@ 4 Concavity powers of strictly log-concave measures
We now explain how the quantity in (48) is connected to spectral inequalities. Fix a smooth and strictly convex $K\in\mathcal{K}^{n}$, and let $L\in\mathcal{K}^{n}$ be smooth and strictly convex as well. Define the function $f:\mathbb{S}^{n-1}\to\mathbb{R}$ by
$f=h_{L}-h_{K},$ (49)
@@
$p(\mu,K)=\sup_{p\in(0,\infty)}\left\{p:\left.\frac{d^{2}}{dt^{2}}\mu(K(f,t))^{p}\right|_{t=0}\leq 0\quad\forall f\in C^{1}(\mathbb{S}^{n-1}),\text{ where }K(f,t)=[h_{K}+tf]\right\},$
where we have that linear combinations of differences of support functions in $C^{2}(\mathbb{S}^{n-1})$ are dense in $C^{1}(\mathbb{S}^{n-1})$. With the notation $\rho=f(\nu_{K})$ (which is in $C^{1}(\partial K)$ by the regularity of $K$), if we now write
@@
In the case where the convex sets $K$ and $L$ are origin symmetric, the admissible perturbations $f$ in (49) must be even. Consequently, the infimum in (50) is taken over all even functions in $C^{1}(\partial K)$. Using the $L_{2}$ method, the functional $J(\rho)$ can be estimated from below using only integrals over $K$; see, for example, Proposition 3.16 in [30]. This has been the prevailing approach in all prior results in the literature.
@@
The main result of this section (Theorem 1.3 in the introduction) reads as follows. We recall that for a vector field $X\in C^{1}(\partial K,\mathbb{R}^{n})$, $\nabla_{\partial K}\cdot X$ is the tangential divergence of $X$.
@@
for every $X\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{n})$. With this notation, we define on $H^{1}(\partial K,\mu)$ the operator $\mathcal{L}$ in Theorem 4.3 by
@@
Thus, the function $\overline{\rho}$ is a weak solution of $\mathcal{L}(\overline{\rho})=1$. Notice that, if $K$ is smooth, then the operator $\mathcal{L}$ is uniformly elliptic since ${\rm II}^{-1}$ is positive definite. Moreover, for every $\rho\in H^{1}(\partial K,\mu)$ a quick computation shows that
$\int_{\partial K}\rho\mathcal{L}(\rho)\,\mathrm{d}\mu=\langle\rho,\rho\rangle_{\rm P}.$ (64)